Shift report 20000722 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 22 2000 - 08:06:44 EDT

  • Next message: Peter H. L. Christiansen: "Good data"

    ------------------------------------------------------------
    Date: Sat Jul 22 08:06:43 2000, Shift: 00:00-08:00
    Supervisor: FR, with: ZIS
    Log book#3 pages 43 - 44
    
    Shift Summary:
    ------------------------------------------------------------
    -RHIC had quench interlok problem and power supply failure.
     Beam was restored at 4:15 but all attemps to ramp were
     unsuccessful. Current status "Ramp tuning continues".
    
    -Runs 1885 and 1886 are pedestal runs.
    
    -There was a problem with the MUMF5 mainframe (Tiles) on 
     BRAHVO. Status was on error and all channels were off. 
     The problem was fixed with the help of Kostantin (we
     just turned it down and on again.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 22 2000 - 08:10:20 EDT